# RDD: Robust Feature Detector and Descriptor using Deformable Transformer

### v4

@@ 7.1 MegaDepth-View
MegaDepth-View is derived from the test scenes of MegaDepth [22]. MegaDepth [22] is a large-scale outdoor dataset containing over 1 million Internet images from 196 different locations. Camera poses are reconstructed using COLMAP [40, 41], and depth maps are generated via multi-view stereo. The test scenes comprise 8 distinct locations around the world, ensuring diversity in the testing data.
@@ 7.2.1 Data Collection
3D reconstruction from imagery captured at multiple altitudes has increasingly garnered attention, driven by the growing UAV industry. Finding reliable correspondences between cross-view imagery has become a significant bottleneck in this domain [4]. RDD is designed to maintain robustness under large camera baselines and aims to enhance the accuracy and reliability of its downstream applications like 3D reconstruction for cross-view imagery. To validate RDD’s ability to address such bottlenecks and evaluate its robustness, we collected the first large benchmark dataset focusing on cross-view imagery. This dataset includes a total of 41 famous locations around the world, such as the Eiffel Tower, Louvre Museum, Sacré-Cœur Basilica, London Bridge, Ponte Vecchio, Las Vegas Strip, Altare della Patria, Flatiron Building, Jackson Square, and Plaza de España. The dataset has around 27,000 images and over 600,000 air-to-ground image pairs.
Inspired by MegaDepth [22], we use COLMAP [40, 41] to reconstruct camera poses and estimate depth maps. Differing from MegaDepth [22] which uses internet images, we collect Internet drone videos and ground images. Drone videos allow us to track frames from the ground up to the air, generating one comprehensive 3D reconstruction including both ground images and frames extracted from drone videos.
@@
Method @$5^{\circ}$$10^{\circ}$$20^{\circ}$
@@
DKM [9] CVPR’23 65.0 77.2 85.7
RoMa [12] CVPR’24 71.3 82.4 89.5
@@
LoFTR [42] CVPR’21 21.5 33.8 45.6
ASpanFormer [5] ECCV’22 45.8 60.0 71.0
ELoFTR [47] CVPR’24 49.4 62.8 73.2
XFeat* [33] CVPR’24 12.0 19.2 27.9
@@
SP [6]+SG [38] CVPR’19 42.0 56.3 67.7
SP [6]+LG [24] ICCV’23 47.9 62.7 73.9
RDD+LG [24] ICCV’23 55.1 68.9 78.9
@@
Figure 7: Qualitative Results on Air-to-Ground. RDD and RDD* demonstrate the ability to extract robust descriptors that perform well in cross-view settings, highlighting the effectiveness of our proposed method. DeDoDe-G* also achieves competitive performance, benefiting from the powerful foundation model Dino-v2 [32].
The raw depth maps obtained from COLMAP often include significant outliers that negatively affect the accuracy of warping used to estimate overlaps and compute matching pixels. These outliers arise primarily from unmatchable moving objects and regions such as the sky or uniform foreground areas like roads. To address these challenges, we implement a series of depth post-processing steps. First, we use a semantic segmentation model [51] to to mask predefined classes prone to unreliable depth estimation, such as sky, sidewalks, vehicles, people, and animals and etc., which are prone to producing unreliable depth information. Second, small and isolated regions are removed using connected component analysis, discarding regions smaller than 1,000 pixels to retain only significant structures. These steps effectively enhance the quality of the depth data by mitigating noise and focusing on stable, meaningful features.
@@ 7.2.2 Results
Recalling the experiment setting in Sec. 4.1 of the main paper, we report the AUC of the recovered pose under threshold (5°, 10° ,and 20°). We use RANSAC to estimate the essential matrix. We compared RDD against the other detector/descriptor methods [6, 44, 49, 50, 13, 10, 33]. The results are presented in Tab. 6 and visually in Fig. 7. Our results show a performance gain compared to previous methods. These results further confirm the robustness of RDD and RDD* in challenging scenarios.
@@ 9 Running Time Analysis
In this section, we present a detailed timing analysis of RDD in both sparse and semi-dense matching settings. We also compare RDD against other methods [6, 44, 49, 50, 13, 10, 33], using the same experimental settings as the main paper. All methods are evaluated on an NVIDIA RTX 4090 GPU with 24GB of VRAM. Tab. 7 shows the inference speeds of all methods, measured in milliseconds. AUC@5°on MegaDepth-1500 for all methods is provided for reference. RDD and RDD* demonstrate competitive feature matching performance with competitive efficiency. A detailed breakdown of the time required for each step of our method is presented in Tab. 8. Notably, RDD is significantly faster than RDD*, as it uses fewer keypoints and does not require refinement. Although RDD* takes more time compared to RDD , it still achieves a good balance between efficiency and performance.
@@
SuperPoint [6] CVPRW’18 302 24.1
DISK [44] NeurIPS’20 98 38.5
ALIKED [50] TIM’23 182 41.8
XFeat [33] CVPR’24 32 24.0
DeDoDe-G [10] 3DV’24 382 47.2
@@ 10 More Experiments
RDD+DeDoDe-G [10] 3DV’24 48.6 65.3 78.0 44.1 58.8 70.3
SuperPoint [6]+RDD CVPRW’18 30.3 48.7 66.5 33.3 51.3 66.9
SuperPoint [6] CVPRW’18 24.1 40.0 54.7 7.50 13.3 21.1
DeDoDe-V2-G [11, 10] CVPRW’24, 3DV’24 47.2 63.9 77.5 33.1 47.6 60.2

### v5

@@ 7.1 MegaDepth-View
MegaDepth-View is derived from the test scenes of MegaDepth [li2018megadepth]. MegaDepth [li2018megadepth] is a large-scale outdoor dataset containing over 1 million Internet images from 196 different locations. Camera poses are reconstructed using COLMAP [schoenberger2016sfm, schoenberger2016mvs], and depth maps are generated via multi-view stereo. The test scenes comprise 8 distinct locations around the world, ensuring diversity in the testing data.
@@ 7.2.1 Data Collection
3D reconstruction from imagery captured at multiple altitudes has increasingly garnered attention, driven by the growing UAV industry. Finding reliable correspondences between cross-view imagery has become a significant bottleneck in this domain [chen2024geometryawarefeaturematchinglargescale]. RDD is designed to maintain robustness under large camera baselines and aims to enhance the accuracy and reliability of its downstream applications like 3D reconstruction for cross-view imagery. To validate RDD’s ability to address such bottlenecks and evaluate its robustness, we collected the first large benchmark dataset focusing on cross-view imagery. This dataset includes a total of 41 famous locations around the world, such as the Eiffel Tower, Louvre Museum, Sacré-Cœur Basilica, London Bridge, Ponte Vecchio, Las Vegas Strip, Altare della Patria, Flatiron Building, Jackson Square, and Plaza de España. The dataset has around 27,000 images and over 600,000 air-to-ground image pairs.
Inspired by MegaDepth [li2018megadepth], we use COLMAP [schoenberger2016sfm, schoenberger2016mvs] to reconstruct camera poses and estimate depth maps. Differing from MegaDepth [li2018megadepth] which uses internet images, we collect Internet drone videos and ground images. Drone videos allow us to track frames from the ground up to the air, generating one comprehensive 3D reconstruction including both ground images and frames extracted from drone videos.
@@
Method $5^{\circ} @$$10^{\circ} @$$20^{\circ}$
@@
DKM [edstedt2023dkm] CVPR’23 65.0 77.2 85.7
RoMa [edstedt2024roma] CVPR’24 71.3 82.4 89.5
@@
LoFTR [sun2021loftr] CVPR’21 21.5 33.8 45.6
ASpanFormer [chen2022aspanformer] ECCV’22 45.8 60.0 71.0
ELoFTR [wang2024eloftr] CVPR’24 49.4 62.8 73.2
XFeat* [potje2024cvpr] CVPR’24 12.0 19.2 27.9
@@
SP [superpoint2018]+SG [sarlin20superglue] CVPR’19 42.0 56.3 67.7
SP [superpoint2018]+LG [lindenberger2023lightglue] ICCV’23 47.9 62.7 73.9
RDD+LG [lindenberger2023lightglue] ICCV’23 55.1 68.9 78.9
@@
Figure 7: Qualitative Results on Air-to-Ground. RDD and RDD* demonstrate the ability to extract robust descriptors that perform well in cross-view settings, highlighting the effectiveness of our proposed method. DeDoDe-G* also achieves competitive performance, benefiting from the powerful foundation model Dino-v2 [oquab2024dinov2].
The raw depth maps obtained from COLMAP often include significant outliers that negatively affect the accuracy of warping used to estimate overlaps and compute matching pixels. These outliers arise primarily from unmatchable moving objects and regions such as the sky or uniform foreground areas like roads. To address these challenges, we implement a series of depth post-processing steps. First, we use a semantic segmentation model [zhou2017scene] to to mask predefined classes prone to unreliable depth estimation, such as sky, sidewalks, vehicles, people, and animals and etc., which are prone to producing unreliable depth information. Second, small and isolated regions are removed using connected component analysis, discarding regions smaller than 1,000 pixels to retain only significant structures. These steps effectively enhance the quality of the depth data by mitigating noise and focusing on stable, meaningful features.
@@ 7.2.2 Results
Recalling the experiment setting in Sec. 4.1 of the main paper, we report the AUC of the recovered pose under threshold (5°, 10° ,and 20°). We use RANSAC to estimate the essential matrix. We compared RDD against the other detector/descriptor methods [superpoint2018, tyszkiewicz2020disk, Zhao2022ALIKE, zhao2023aliked, gleize2023silksimplelearned, edstedt2024dedode, potje2024cvpr]. The results are presented in Tab. 6 and visually in Fig. 7. Our results show a performance gain compared to previous methods. These results further confirm the robustness of RDD and RDD* in challenging scenarios.
@@ 9 Running Time Analysis
In this section, we present a detailed timing analysis of RDD in both sparse and semi-dense matching settings. We also compare RDD against other methods [superpoint2018, tyszkiewicz2020disk, Zhao2022ALIKE, zhao2023aliked, gleize2023silksimplelearned, edstedt2024dedode, potje2024cvpr], using the same experimental settings as the main paper. All methods are evaluated on an NVIDIA RTX 4090 GPU with 24GB of VRAM. Tab. 7 shows the inference speeds of all methods, measured in milliseconds. AUC@5°on MegaDepth-1500 for all methods is provided for reference. RDD and RDD* demonstrate competitive feature matching performance with competitive efficiency. A detailed breakdown of the time required for each step of our method is presented in Tab. 8. Notably, RDD is significantly faster than RDD*, as it uses fewer keypoints and does not require refinement. Although RDD* takes more time compared to RDD , it still achieves a good balance between efficiency and performance.
@@
SuperPoint [superpoint2018] CVPRW’18 302 24.1
DISK [tyszkiewicz2020disk] NeurIPS’20 98 38.5
ALIKED [zhao2023aliked] TIM’23 182 41.8
XFeat [potje2024cvpr] CVPR’24 32 24.0
DeDoDe-G [edstedt2024dedode] 3DV’24 382 47.2
@@ 10 More Experiments
RDD+DeDoDe-G [edstedt2024dedode] 3DV’24 48.6 65.3 78.0 44.1 58.8 70.3
SuperPoint [superpoint2018]+RDD CVPRW’18 30.3 48.7 66.5 33.3 51.3 66.9
SuperPoint [superpoint2018] CVPRW’18 24.1 40.0 54.7 7.50 13.3 21.1
DeDoDe-V2-G [edstedt2024dedodev2, edstedt2024dedode] CVPRW’24, 3DV’24 47.2 63.9 77.5 33.1 47.6 60.2